Hydrogen is prepared commercially by the reaction of methane and water vapor at elevated temperatures.

CH4 (g) + H2O (g) ⇌ H2 (g) + CO (g)

What is the equilibrium constant for the reaction if a mixture at equilibrium contains gases with the following concentrations, at a temperature of 760 °C?

CH4 = 0.126 M

H2O = 0.242 M

CO = 0.126 M

H2 = 1.15 M



1
Expert's answer
2022-01-18T04:00:04-0500

[CH4] = 0.126 M

[H2O] = 0.242 M

[CO] = 0.126 M

[H2] = 1.15 M


Solution:

Balanced chemical equation:

CH4(g) + H2O(g) ⇌ 3H2(g) + CO(g)

The equilibrium expression for the reaction system:

Kc = ([CO] × [H2]3) / ([H2O] × [CH4])

Kc = (0.126 × 1.153) / (0.242 × 0.126) = (0.19163) / (0.030492) = 6.28

Kc = 6.28 M2


Answer: The equilibrium constant (Kc) for the reaction is 6.28